The nurse is preparing potassium bicarbonate (TriLyte) for a pediatric client scheduled for a colonoscopy. The directions state to provide the prep at an estimated rate of 25 mL/kg/hr. The child weighs 55 lb. The nurse should encourage the child to drink approximately how much of the solution each hour? (round to the nearest 1/2 a cup.)

3 cups

2 cups

3.5 cups

1 cup

2.5 cups
Answer #1

The weight of child is 55 lb. So first we will convert weight into Kg

Since 1 Kg = 2.2 lb

So, 55 lb = 55/2.2 Kg

=25 Kg

So child weight is 25 Kg.

The direction state is to provide the preparation at an estimated rate of 25 mL/kg/hr.

So for 25 Kg body weight, the dose administered every hour will be calculated as

25\times 25 = 625 mL/hr

Now, One cup is equal to just under 236.6 milliliters, but in nutrition labelling, one cup is equal to 240 milliliters.

For 625 mL, the number of cups will be

So, 625/240 = 2.6 cup

So, to round nearest 1/2 a cup, the child has to drink 2.5 cup.